OpenStudy (anonymous):

how to solve: lim (x-->0) sin(2x)sin(3x)/sin(3x^2)

OpenStudy (shayaan_mustafa):

it is look like this \[\large \lim_{x \rightarrow 0}\frac{\sin(2x)\sin(3x)}{\sin(3x ^{2})}\]Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

We know that \[\lim_{x\to 0} \frac{ \sin x}{x}=1-------------------(1)\] and also \[\lim_{x\to 0} \frac{ \sin x^2}{x^2}=1-------------------(2)\] so her we have \[\lim_{x\to 0} \frac{ \sin 2x\sin 3x }{\sin 3x^2 }\] We have to create forms like (1) and (2) to evaluate the limit Dividing and mulitplying by \( 2x \times 3x\) \[\lim_{x\to 0} \frac{ \sin 2x\sin 3x \times 2x \times 3x }{\sin 3x^2 \times 2x \times 3x }=1\] So we have now \[\lim_{x\to 0} \frac{\sin 2x }{2x} \times \frac{ \sin 3x}{3x} \times \frac{ 2x \times 3x}{ \sin 3x^2}\] \[\lim_{x\to 0} \cancel {\frac{\sin 2x }{2x}} 1 \times \cancel{\frac{ \sin 3x}{3x}} 1 \times 2\frac{ 3x^2}{ \sin 3x^2}\] Now we have \[\lim _{x\to 0} 2\cancel{\frac{ 3x^2}{ \sin 3x^2}}1\] so we get \[\lim_{x\to 0} \frac{ \sin 2x\sin 3x }{\sin 3x^2 }=2\]

OpenStudy (anonymous):

thank you!

OpenStudy (ash2326):

@azalea Welcome to Open Study:)